OpenStudy (anonymous):

Simplify: ( y−7 × y−3 )−1

OpenStudy (anonymous):

\((y-7 \times y-3)-1\) Is this your question?

OpenStudy (anonymous):

ya

OpenStudy (anonymous):

the number are in the top

OpenStudy (anonymous):

You mean that the numbers are exponents?

OpenStudy (anonymous):

yaa

OpenStudy (anonymous):

\((y^{-7} \times y^{-3})^{-1}\)

OpenStudy (anonymous):

yess

OpenStudy (anonymous):

Raising a power to a power - multiply the exponents. \((x^2)^3 = x^6\)

OpenStudy (cggurumanjunath):